A weather balloon near sea level (101,377 Pa, 298.56 K) has a volume of 270.83 ft3 (cubed). If it is launched and ascends to a height of 105,200 feet (857.06 Pa, 227.85 K) before bursting. What was its final volume, in feet cubed?
Answer #1

Final volume = 24448 ft3

Explanation

According to ideal gas law,

(P1 * V1) / (T1) = (P2 * V2) / (T2)

where P1 = initial pressure = 101377 Pa

V1 = initial volume = 270.83 ft3

T1 = initial temperature = 298.56 K

P2 = final pressure = 857.06 Pa

V2 = final volume

T2 = final temperature = 227.85 K

Substituting the values,

[(101377 Pa) * (270.83 ft3)] / (298.56 K) = [(857.06 Pa) * (V2)] / (227.85 K)

V2 = (270.83 ft3) * (101377 Pa / 857.06 Pa) * (227.85 K / 298.56 K)

V2 = (270.83 ft3) * (118.2846) * (0.76316)

V2 = 24447.95 ft3